Cleaned floor and began a 10cm cut in meter 53S161E. It became quickly apparent that there was a soil change about 3cm below the surface at approximately 71cm b.d. Test cuts with a trowel of 3-5cm in meters 51S161E and 52S161E were done to determine if the soil change was uniform across locus 6, but at first the uniformity did not seem to exist. The test cut in meter 51S161E revealed the soil to be the same as in locus 6, but the test cut in meter 52S161E revealed a soil to be similar to the soil in locus 4. Depths were then taken to see if that may influence the discrepancy in soil types. Meter 51S161E had a depth of approximately 64cm b.d. It was decided to try and level the trench floor.
3-5cm cuts were done in meters 51S160.5E, 51S161E, 52S160.5E, 52S161E, bringing the depths down to the respective depths of 66, 68, 66, 66cm b.d. Pictures will be taken tomorrow and mag duties are tomorrow
so locus determinations will be made on thursday. 1/8 cassetta box of terracotta and 1/2 latte box of bulk pottery were found from locus 6.
|Property or Relation||Value(s)|
[Standard: Dublin Core Terms]
Open Context editors work with data contributors to annotate datasets to shared vocabularies, ontologies, and other standards using 'Linked Open Data' (LOD) methods.
The annotations presented above approximate some of the meaning in this contributed data record to concepts defined in shared standards. These annotations are provided to help make datasets easier to understand and use with other datasets.
To the extent to which copyright applies, this content
carries the above license. Follow the link to understand specific permissions
Required Attribution: Citation and reference of URIs (hyperlinks)